# Machine learning of the Ising model on a spherical Fibonacci lattice

### v2

@@ IV Antiferromagnetic Ising model on a spherical Fibonacci lattice
We now consider the case of negative spin coupling, $J<0$. In this scenario, the nearest spins tend to anti-align, leading to the formation of an ordered antiferromagnetic phase. Consequently, the connectivity features of the lattice will influence the ground-state structures. We set $J=-1$ and focus on the spherical Fibonacci lattice with $N=1000$ and $r_{c}/R=0.1298$. At zero temperature, the MC simulation predicts that most neighboring spins are oppositely aligned. However, the inhomogeneity of the spherical lattice leads to some nearest neighbors exhibiting parallel alignment. The details of the ground-state spin configuration are presented in the left panel of the top row of Fig.12. As the temperature increases, while remaining low enough, the spin distribution stays in an ordered state, though it becomes irregular at certain sites. We present the corresponding spin configuration at $T/|J|=2.0$ as viewed from two different directions in the middle and right panels of the top row of Fig.12. As the temperature increases sufficiently, the ordered structure is gradually disrupted, resulting in a disordered state similar to that observed in the ferromagnetic Ising model (shown in the bottom row of Fig.3).
@@
For 2D square lattice, the phase transition temperature of the antiferromagnetic Ising model can also be theoretically predicted, $T^{\Box}_{c}/J\simeq 2.269$[31], which is the same as that of the ferromagnetic Ising model. To determine $T_{c}$ for the spherical Fibonacci lattice, we will continue applying the previous methods. However, due to the less regular spin patterns and the presence of irregular neighbors on the sphere, the antiferromagnetic model exhibits exotic boundaries even at zero temperature. Therefore, the previous methods for estimating entropy and magnetic susceptibility are not applicable in this case. We can only calculate the specific heat by using $C_{V}=\left(\frac{\partial E}{\partial T}\right)_{V}$ and Eq.(3).
@@
Figure 13: (Top panel) Total energy as a function of temperature for the spherical antiferromagnetic Ising model with $N=1000$. (Middle panel) Specific heat as a function of temperature for the same model. (Bottom panel) Specific heat as a function of temperature, calculated using Eq.(3).
Following the same procedure, we compute the total energy for stable spin configurations at each temperature and plot $E$ vs $T$ in Fig.13 (top panel). By fitting the $E-T$ curve with spine functions and taking derivatives with respect to $T$, we obtain the specific heat $C_{V}$, which is also presented in Fig.13 (middle panel). Interestingly, the total energy near zero temperature is not the lowest, which contrasts with the behavior observed in the planar square lattice. This can be explained as follows. The planar square lattice is a bipartite lattice that can be divided into two sub-lattices, with sites from one sub-lattice interacting only with those from the other. If the spins on the two sub-lattices have opposite signs, the total energy is minimized. In the case of the Fibonacci lattice, the sites with 3 or 5 neighbors break this symmetry, leading to an increase in total energy, so that the most uniform spin distribution does not correspond to the lowest energy state. This is, in fact, a distinctive feature of non-uniform lattices. Furthermore, based on the behavior of $C_{V}$, we obtain an estimation of the critical temperature: $T_{c}/|J|=2.07$, which is relatively lower than that of the ferromagnetic case. This is understandable. In an ordered phase at very low temperatures, most sites have 4 neighbors, of which 2 share the same spin. As the temperature increases, it costs $4J$ for such a site to flip its spin, which is only half of the cost in the ferromagnetic Ising model. Consequently, the phase transition temperature is effectively lowered. After performing the Bootstrap error analysis, the result is refined to $T_{c}=2.059\pm 0.016J$.
To cross-check our results, we also employ Eq.(3) to analyze the behavior of $C_{V}$ and present the numerical results in the bottom panel of Fig.13. Notably, the oscillations in this case are more pronounced than those in the ferromagnetic scenario. This is because the spin pattern of the ordered phase is less regular compared to the latter. In the ferromagnetic case, the ordered spin configuration is relatively simple, with most sites sharing the same spin orientation. By fitting the data with a Gaussian function, we estimate the critical temperature as $T_{c}=1.95J$. Furthermore, Bootstrap error analysis refines this estimate to $T_{c}=1.932\pm 0.167J$.
@@
Finally, we apply the method of GCN to estimate the phase transition temperature. The corresponding numerical results are visualized in the top panel of Fig.14, where the critical temperature is found to be $T_{c}/|J|=2.46$. In the bottom panel, we present the linear fittings of $p_{\text{o}}(T)$ and $p_{\text{d}}(T)$ near $T_{c}$. Using Eq. (5), the estimated uncertainty is $T_{c}=2.4895\pm 0.0610J$. The value of $T_{c}$ is fairly close to that obtained for the ferromagnetic case using the same method. This similarity may arise from the resemblance between the spin configurations of the two models at high temperatures, where both exhibit similarly chaotic spin distributions. Furthermore, the estimated $T_{c}$ is also relatively higher than that derived from the behavior of $C_{V}$. This discrepancy could be due to the fact that, in the antiferromagnetic case, the distinction between the ordered and disordered phases near $T_{c}$ is less pronounced than in the ferromagnetic case, which affects the GCN’s ability to differentiate between the two phases. Therefore, we believe that the $T_{c}$ obtained from the former method is relatively more accurate. Additionally, $p_{\text{o}}$ and $p_{\text{d}}$ exhibit non-smooth behavior in the high-temperature regime. This irregularity arises from the inherent ambiguity in the similarity between spin configurations at low and high temperatures in the spherical antiferromagnetic Ising model, leading to suboptimal performance of the GCN in classification tasks. Nevertheless, despite these challenges, the GCN-based approach remains a valuable tool for identifying the phase transition temperature of the model.
@@ V Conclusion
In this paper, we studied the Ising model on a spherical Fibonacci lattice, a structure that balances relative uniformity with irregular sites that significantly influence spin behavior. Employing Monte Carlo simulations and graph convolutional networks (GCNs), we analyzed spin configurations and identified phase transition temperatures for both ferromagnetic and antiferromagnetic cases.
In the ferromagnetic scenario, sites with fewer neighbors were more prone to spin flips at low temperatures, an effect that weakened as temperature rose. This behavior led to a phase transition temperature higher than that of the planar Ising model, highlighting how the lattice’s curvature and connectivity bolster ferromagnetic order. Conversely, in the antiferromagnetic case, lattice irregularities prevented the total energy from reaching its minimum at zero temperature-unlike an ideal bipartite lattice-revealing the complex interplay between geometry and spin interactions in non-planar systems.
By utilizing specific heat analysis and GCNs, which adeptly capture the lattice’s geometric properties, we accurately estimated phase transition temperatures for both cases. Notably, finite size effects were minimal within the system sizes studied, likely due to the uniform distribution of sites in the spherical Fibonacci lattice, mitigating size-dependent variations common in other lattice types. These findings underscore the pivotal role of geometric features in shaping spin dynamics and phase transitions, distinguishing this system from traditional planar models.
@@ VI Acknowledgments
H.G. was supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904) and the National Natural Science Foundation of China (Grant No. 12074064). X. Y. H. was supported by the the National Natural Science Foundation of China (Grant No. 12405008) and the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611).
@@ Appendix A Bootstrap Error Analysis Method
    • Confidence Interval: Take the 2.5% and 97.5% percentiles of $\{T_{c}^{(b)}\}$ to define the 95% confidence interval $[T_{c}^{\text{low}},T_{c}^{\text{high}}]$.
Appendix B Method of estimating the entropy
Figure 15: Boundary of a stable spin configuration and its zoom-in view.
To calculate the entropy of the spherical Ising model, we employ the Boltzmann formula:
$S=k_{B}\ln\Omega,$ (6)
where the Boltzmann constant is set to $k_{B}=1$ in natural unites. For a model with $N_{c}$ “proper” sites at the boundaries, the total number of possible states is given by $\Omega=2^{N_{c}}$ since each spin has two orientations. This yields $S=N_{c}\ln 2$. Thus, our primary focus is on determining the “proper” number $N_{c}$ of boundary sites .
Figure 16: A special local structure of the spherical lattice.
The left panel of Fig.15 presents a stable spin configuration, and a zoomed-in view of a part of its boundary is shown in the right panel. Here the boundary refers to the interface between regions of opposite spin orientations, visually represented by the white and black lattice sites. To illustrate the method for determining $N_{c}$, we select a lattice site and identify its nearest neighbor with an opposite spin. For instance, the nearest neighbor of lattice site 2 with an opposite spin is lattice site $B$, and so forth. However, this method results in double-counting of the boundary lattice sites, yielding $N_{c}=6+6=12$. To correct for this, we include the lattice sites themselves in the counting process. Specifically, for lattice site 2, we record its nearest neighbor with opposite spin as site $B$, denoted as $(2,B)$. Similarly, for site 3, we record $(3,C)$, and so on. Likewise, for lattice site $B$, we record $(B,2)$. Since $(2,B)$ and $(B,2)$ are identical, we eliminate one duplicate entry. This refinement leads to a final determination of $N_{c}=6$.
Now consider a more complex configuration. For instance, Fig.16 illustrates the local lattice structure of the spherical Ising model, where a single spin at site 7 is flipped downward. The boundary of this site consists of four neighboring lattice sites $(a,b,c,d)$. To ensure consistency with the condition for determining shown in Fig.15, we also include the lattice site itself as part of the boundary. Consequently, in this case, the ”proper” number of boundary sites is 5: $(7,a,b,c,d)$, leading to $N_{c}=5$.
Appendix C Methodology of phase classification: Details of GCN

### v3

@@ IV Antiferromagnetic Ising model on a spherical Fibonacci lattice
We now consider the case of negative spin coupling, $J<0$. In this scenario, the nearest spins tend to anti-align, leading to the formation of an ordered antiferromagnetic phase. Consequently, the connectivity features of the lattice will influence the ground-state structures. We set J=1J=-1 and focus on the spherical Fibonacci lattice with $N=1000$ and $r_{c}/R=0.1298$.
In contrast to the ferromagnetic case on the spherical Fibonacci lattice, Monte Carlo simulations at zero temperature reveal that while most neighboring spins align antiparallel, the nonuniform geometry of the spherical lattice, where certain sites have 3 or 5 neighbors, prevents all local interactions from being simultaneously satisfied. As a result, some neighboring spins remain aligned in the same direction. One of the ground-state configuration is shown in the upper-left panel of Fig. 12. This behavior exemplifies the well-known phenomenon of geometric frustration. Geometric frustration refers to situations where, due to geometric constraints of the lattice, local interactions cannot all be simultaneously optimized, preventing the system from establishing a perfectly ordered state Moessner and Ramirez (2006); Ronceray and Le Floch (2019). A classic example is the antiferromagnetic triangular lattice: placing one spin at each vertex of an equilateral triangle with antiferromagnetic interactions, two spins can align antiparallel to satisfy their interaction, but the third spin cannot simultaneously satisfy antiferromagnetic interactions with both neighbors. As a consequence, geometric frustration leads to the existence of a large number of degenerate ground states at zero temperature. In other words, even when the system reaches its lowest energy, it cannot establish a unique globally ordered configuration. As the temperature increases, while remaining low enough, the spin distribution stays in an ordered state, though it becomes irregular at certain sites. We present the corresponding spin configuration at $T/|J|=2.0$ as viewed from two different directions in the middle and right panels of the top row of Fig.12. As the temperature increases sufficiently, the ordered structure is gradually disrupted, resulting in a disordered state similar to that observed in the ferromagnetic Ising model (shown in the bottom row of Fig.3).
@@
For 2D square lattice, the phase transition temperature of the antiferromagnetic Ising model can also be theoretically predicted, $T^{\Box}_{c}/J\simeq 2.269$Wadati and Akutsu (1988), which is the same as that of the ferromagnetic Ising model. To determine $T_{c}$ for the spherical Fibonacci lattice, we will continue applying the previous methods. However, due to the less regular spin patterns and the presence of irregular neighbors on the sphere, the antiferromagnetic model exhibits exotic boundaries even at zero temperature. Therefore, the previous methods for estimating magnetic susceptibility are not applicable in this case. We can only calculate the specific heat by using $C_{V}=\left(\frac{\partial E}{\partial T}\right)_{V}$ and Eq.(5).
@@
Figure 13: (Top panel) Total energy as a function of temperature for the spherical antiferromagnetic Ising model with $N=1000$. (Middle panel) Specific heat as a function of temperature for the same model. (Bottom panel) Specific heat as a function of temperature, calculated using Eq.(5). The black dashed lines represent the raw Monte Carlo data points, and the red solid line shows the Gaussian fit around the peak.
Following the same procedure, we compute the total energy for stable spin configurations at each temperature and plot $E$ vs $T$ in Fig.13 (top panel). By fitting the ETE-T curve with spline functions and taking derivatives with respect to $T$, we obtain the specific heat $C_{V}$, which is also presented in Fig.13 (middle panel).
Based on the behavior of $C_{V}$, we obtain an estimation of the critical temperature: $T_{c}/|J|=2.00$, which is relatively lower than that of the antiferromagnetic square lattice. This is understandable. As previously mentioned, due to the nonuniform distribution of sites on the spherical lattice, some lattice points have an odd number of neighbors (3 or 5), leading to the phenomenon of geometric frustration. The geometric frustration reduces the system’s tendency to establish global order, thereby lowering the phase transition temperature. After performing the Bootstrap error analysis, the result is refined to $T_{c}/|J|=1.999\pm 0.006$. To cross check our results, we also employ Eq.(5) to analyze the behavior of $C_{V}$ and present the numerical results in the bottom panel of Fig.13. By fitting the data with a Gaussian function, we estimate the critical temperature to be $T_{c}/|J|=1.98$. Furthermore, Bootstrap error analysis refines this estimate to $T_{c}/|J|=1.982\pm 0.012$.
@@
Similar to the spherical ferromagnetic model, we estimated the phase transition temperature of the spherical antiferromagnetic Ising model using both GCNs and unsupervised learning approaches. The numerical results from the GCNs are shown in the top panel of Fig.14, yielding a critical temperature of $T_{c}/|J|=1.49$. In the bottom panel, we present the linear fittings of $p_{\text{o}}(T)$ and $p_{\text{d}}(T)$ near $T_{c}$, with the estimated uncertainty calculated via Eq.(7) as $T_{c}/|J|=1.498\pm 0.210$. The results from unsupervised learning are shown in Fig.15: the left panel depicts the clustering outcome, which is unlike the spherical ferromagnetic model and does not show clearly separated clusters; the right panel shows the cluster fractions as a function of temperature, from which the phase transition temperature is estimated to be $T_{c}/|J|=1.806$.
The estimated $T_{c}$ is significantly lower than the value obtained from the specific heat $C_{V}$. This discrepancy can be attributed to two main factors. First, in the antiferromagnetic case, the distinction between the ordered and disordered phases near $T_{c}$ is less obvious than in the ferromagnetic case, which affects the GCNs’ ability to differentiate the two phases, and the unsupervised learning also fails to produce well-separated clusters. Second, supervised learning requires pre-input of the zero-temperature ordered spin configurations; however, the spherical antiferromagnetic Ising model exhibits geometric frustration, leading to a highly degenerate ground state. As a result, the initial training samples do not necessarily cover all possible degenerate ground states, which reduce the sensitivity of the machine learning approach. This discrepancy also reflects a limitation of the machine learning approach: while it performs well for systems with clear phase boundaries (such as the ferromagnetic case), its estimates of $T_{c}$ in frustrated antiferromagnetic systems are less reliable. Even unsupervised learning struggles to generate well-separated clusters, which further indicates that geometric frustration reduces the ability of machine learning to capture the underlying phase transition. Therefore, we consider the $T_{c}$ obtained from the specific heat method to be more reliable.
Figure 15: (Left panel) The similarity of different spin configurations with respect to the two principal components. (Right panel)The cluster fraction, $p_{\text{o}}$ and $p_{\text{d}}$, versus temperature of the antiferromagnetic Ising model on a spherical Fibonacci lattice for $N=1000$ .
@@ V Conclusion
In this paper, we studied the Ising model on a spherical Fibonacci lattice, a structure that balances relative uniformity with irregular sites that significantly influence spin behavior. Employing Monte Carlo simulations, principal component analysis (PCA) and graph convolutional networks (GCNs), we analyzed spin configurations and identified phase transition temperatures for both ferromagnetic and antiferromagnetic cases.
In the ferromagnetic scenario, sites with fewer neighbors were more prone to spin flips at low temperatures, an effect that weakened as temperature rose. This behavior led to a phase transition temperature higher than that of the planar Ising model, highlighting how the lattice’s curvature and connectivity bolster ferromagnetic order. Conversely, in the antiferromagnetic case, lattice irregularities induces geometric frustration, which reveals the complex interplay between geometry and spin interactions in non-planar systems. These findings underscore the pivotal role of geometric features in shaping spin dynamics and phase transitions, distinguishing this system from traditional planar models.
@@
While our study focuses on the Ising and XY models on spherical Fibonacci lattices, the GCNs is broadly applicable to other statistical models and non-Euclidean geometries. For instance, multi-state systems such as the $q$-state Potts model can be naturally incorporated by extending the binary classification task (ordered vs. disordered) to a multi-class task (distinguishing $q$ states), allowing GCNs to effectively identify complex phase diagrams across multiple competing states. Prior to performing the multi-class task, unsupervised learning methods can be used to extractfeatures of spin configurations to better distinguish between different phases.
Beyond spherical lattices, GCNs are well suited for systems defined on other curved geometries. For example,geodesic domes whose triangular mesh structure locally resembles that of Fibonacci lattices, while toroidal embeddings introduce periodic boundary conditions that can be encoded in the adjacency matrix. These features enable the GCNs to capture long-range correlations and topological constraints intrinsic to the geometry.
Finally, combining GCNs with unsupervised learning techniques such as PCA provides a promising approach to uncover intrinsic phase structures in complex systems, helping to reveal subtle correlations that may be difficult to detect with purely supervised learning. Overall, these extensions demonstrate the versatility of the GCNs framework in studying a wide range of statistical systems on non-Euclidean geometries.
@@ VI Acknowledgments
H.G. was supported by the Innovation Program for Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2021ZD0301904) and the National Natural Science Foundation of China (Grant No. 12074064). X. Y. H. was supported by the the National Natural Science Foundation of China (Grant No. 12405008) and the Jiangsu Funding Program for Excellent Postdoctoral Talent (Grant No. 2023ZB611).
@@ Appendix A Bootstrap Error Analysis Method
    • Confidence Interval: Take the 2.5% and 97.5% percentiles of $\{T_{c}^{(b)}\}$ to define the 95% confidence interval $T_{c}^{\text{low}},T_{c}^{\text{high}}]$.
Appendix B Methodology of phase classification: Details of GCNs